A block is placed on top of a fixed smooth inclined plane inclined at 30° to horizontal. The length of plane is 5 m. The block slides down the plane and reaches at bottom. The speed of the block at bottom will be nearly


1
Expert's answer
2021-08-03T11:42:49-0400

"\text{By the condition of the problem, }\newline
\text{we have a smooth inclined plane.}"

"\text{That is, frictional forces can be neglected.}"

"l = 5m"

"h = l*\sin30\degree= 2.5m"

"E = E_p+E_k"

"\text{At the top of an inclined plane:}"

"E_p = mgh;E_k=0"

"\text{At the foot of an inclined plane:}"

"E = E_p'+E_k'"

"E_k'=\frac{mv^2}{2};E_p'=0;"


"mgh = \frac{mv^2}{2}"

"v=\sqrt{2gh}=\sqrt{2*9.8*2.5}=7\frac{m}{s}"


"\text{Answer: }v=7\frac{m}{s}"